OpenStudy (mathmath333):

find two-three digit numbers whose sum is multiple of 504 and quotient is multiple of 6..

OpenStudy (mathmath333):

please interpret the que also

ganeshie8 (ganeshie8):

Say the two numbers are \(A\) and \(B\) such that \(A > B\) : \(\large A + B = 504t\) \(\large \dfrac{A}{B} = 6 \implies A = 6B \) combining above both you get : \(\large 6B + B = 504t \implies B = 72t\)

ganeshie8 (ganeshie8):

plugin \(t = 1\), you get \(B = 72*1 = 72\) But you want a 3 digit number, so discard this

ganeshie8 (ganeshie8):

pluging \(t = 2\), whatd do you get ?

OpenStudy (mathmath333):

144

ganeshie8 (ganeshie8):

yes, since it is a 3 digit number you're okay with B how about A ?

ganeshie8 (ganeshie8):

use \(\large \dfrac{A}{B}=6\) to find \(A\)

OpenStudy (mathmath333):

432..?

ganeshie8 (ganeshie8):

yes, t > 2 will not work, why ?

ganeshie8 (ganeshie8):

hey you get B = 144*6 = 864, right ?

OpenStudy (mathmath333):

cuz its 3 digit no